UT 20130327 Technical Observing

Observer: JHill (Tucson)
Telescope Operator: SAllanson (LBTO)
Additional AO Support: JChristou (LBTO)
Telescope Support: JLittle (LBTO)
Instrument Support: TShih (LBTO)


We tested a new version of the LBC software.

We did some mixed mode observations with AGW1(LUCI1) and LBC-Red.

We collected some Step Offset data for Issue 4545.

LBTI worked on closing the AO loop and offsetting through IIF/AOS.


Here's the list of things to do without SX or DX "W" units, without DX AdSec, and without movable DX M3:

  • DONE LBC new software - Kellee has activated the new LBC code with improved diagnostics
  • DONE LBC-Red co-tracking with LUCI1 for Issue 4291
  • DONE PCS test - Michele has started the diagnostic patch version of PCS
  • DONE Issue 4545 Step Offset Data
  • LUCI1-AGW1 focus
  • test BACK_SIZE=32 for MODS1 guiding (already done for LUCI1)
  • LBTI AO and offsets on SX side


02:25 Steve has some issues with dynamic balance as we reconfigure to LBC on right side.


LBC Tests with new software

02:38 Preset to WT10_0252 authorized as LBC on both sides, but LBC is deployed only on the right side. Collimation, pointing and focus are all decent in open-loop. Just a bit of initial coma and focus on Red - easily within capture range.

02:50 Preset to SA98_UsY as an OB test (sky is a bit cloudy) x1 exposures

02:57 Replay with exposure x20 for guiding test

Mixed mode LUCI1 and LBC-Red

These are tests for Issue 4291.

03:05 Authorize LUCI1 @ LFBG and LBC-Red at Right Prime

03:07 SYNC preset from LBC-Red on Right to WT10_0252

03;08 SYNC preset (via IRAF) for LUCI1 to WT10_2052

DX primary runs into X-limit mostly from the 42 arcsec RY and -1.96 mm X pointing correction. Star is at 1109, 2971 on LBC-Red chip2.

Acquire preset on left to center pointing.

03:15 Adjust CA +19.5 on Right to relieve primary limits (a sort of manual range balancing). Leaking is turned off for LBC.

Resend presets - why did the optics move when we did this? Because it rebalanced after our previous adjustment.

03:25 Adjust CA another +30 on right, and IE +17.6 on right.

03:27 Resend presets - now optics are happy. Star is at 1034,2936 on LBC - this is off the rotator center, so the star will orbit the rotator center as we track.


03:38 Fast OB with 100 sec x 20 exposures - LBC-red guided r' images are round and crisp at 0.7 arcsec FWHM

04:00 Fixed the telemetry plots for LBTplot by adding "tcs/" into the path.

04:02 vertical stripes appear on lbcr.20130327.040213.fits

Clouds are increasing (beyond 1 mag), and the 100 sec r' exposures are saturated.

04:30 Turn off LBC

Note that a guideplot centered on 4.6 UT shows binocular guiding (before) and monocular guiding (after).

Issue 4545 Step Offset Data

04:34 ACTIVE preset to BS9143 with LFBG only gs=0 (on-axis)

Doug is producing a script for data collection from home.

I think we are still running near half gain on Z9-Z11=+-0.5, and Z22=-1.0, but the plot shows smooth correction and slow drift for Z11 and Z22. Why? Does that mean the Z22 is predominantly from Z11 cross-talk? Seeing is 0.45 to 0.65 arcsec.

04:51 turn on multi_step_focus.py

04:48 Changing to an off-axis star (still on BS9143) AGW position: -0.0424046, 119.621 (native: -0.224, 492.436)

05:04 SX shell has ripped because of the elevation variable timeout. Active optics has squeezed the image into a 1 arcsec blob. It recovered quickly, AND there was no disturbance in the Z11/Z22 correction (perhaps the rms was always high?).

05:22 Changing to a more off-axis guide star AGW position: -0.00089443, 174.461 (native: -0.357, 437.466)

05:35 Doug testing his new data collection script. First offsetting to center the pupil on an integer lenslet. The scale is 180 arcsec to move the pupil a full subap.

05:47 gs=10 well centered at AGW position: -0.0737014, 174.467 (native: -0.43, 437.46)

05:53 Preset on-axis BS9143 then offset to: AGW position: 0.0393473, 36.0618 (native: 0.124, 576.194)

06:02 Clouds make the wfs star look very faint. This caused Z11 correction on M1 to increase from ~50 to ~220 nm.

06:03 stopWFSing while Doug is collecting data with IDL. 2+ mag of clouds makes occasional interruptions.

06:50 Are the clouds inducing the Z11 changes? And big changes in Z9 and Z20? Maybe we've found the problem.........OR NOT. Could something be influencing the sextractor centroids when the background is high or the star is faint?

06:59 Clear active optics on M1

07:16 Problem with rotator angle in IDL ?? Change to L_FBGRPosition (in arcsec). This was causing the big amounts of Z9, Z10, Z20, Z21 as time passed from the preset.

We plan to continue these measurements (various calibrations related to active optics and Issue 4545) during the second half of tomorrow night.

07:26 Handover to LBTI (Phil, Vidhya, Vanessa)

LBTI AO tests

Vanessa and Alfio at Magellan, Vidhya, Phil and John at 3 places in Tucson.

UBC is under vaccum.

AO loop was able to close with 400 modes (as usual, but now under vacuum).

10:52 multi_step_focus collided with binocular pointing updates when SX preset finally succeeded.

Offset Test (Testing offset pointing commands through IIF)

Must have a live/active preset for these offsets.

12:02 REL offset in DEC on SX failed because Bayside moved in pure Y. RA offset failed similarly a few minutes earlier.

12:05 DETXY offset on SX failed in a similar way

But the Preset was sent in IDLE mode rather than PARALLACTIC mode. Does that also affect the DETXY transformation?

AND LBTI Bayside units are not in millimeters, by x4.

12:29 sent Adaptive ACE_TRACK preset in parallactic mode

Alfio adding scale parameters in wfshard.conf (or some name like that)

12:34 restart wfs arbitrator to ingest those parameters (incorrectly)

12:41 offset 3 0 in DETXY REL ----> AOS says Y=-1.785 mm

12:50 offset 3 0 in DETXY REL ---->AOS says 0, 0

AcquireREF and close loop manually.....

12:54 offset 3 0 in DETXY REL ----> AOS says Y=-1.785 mm

12:55 offset 3 0 RADEC REL --- still moved purely in Y, but parallactic angle is 88 deg

12:56 restart wfs arbitrator to ingest those parameters

13:08 offset 3 0 in DETXY REL now the scale is corrected but we don't know about the sign.


13:10 Close for sunrise

-- JohnHill - 27 Mar 2013

Topic revision: r5 - 29 Mar 2013, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback